When exchanging US Dollars (USD) for Philippine Peso (PHP) the number of Philippine Pesos received is directly proportional to the number of US Dollars to be exchanged. If 1,050 USD can be converted into 46,791.15 PHP.
Find the constant of proportionality k.
k =
(If needed, round answer to 3 decimal places.)
Using the k from above find the amout of PHP given that you have 1,200 USD to convert.
you will receive
PHP (If needed, round answer to 2 decimal places.)

Answer by mananth(16946)

When exchanging US Dollars (USD) for Philippine Peso (PHP) the number of Philippine Pesos received is directly proportional to the number of US Dollars to be exchanged. If 1,050 USD can be converted into 46,791.15 PHP.
Find the constant of proportionality k.
P alpha D
P= k*d
k =P/D
k= 46,791.15/1050
k=44.563
Again P= k*D
P = 44.563*1200
P= 53475.60